# Vector Field-Guided Learning Predictive Control for Motion Planning of Mobile Robots with Unknown Dynamics

### v1

@@ V Theoretic Analysis
This section presents the theoretical analyses regarding Theorem 1. We prove that the Eb-RHRL algorithm, integrating the game-based barrier function, converges to the optimal value function and control, thereby solving the OMP problem.
The convergence of the finite-horizon adaptive critic learning algorithms has been proved in [41] for deterministic systems, and the work [40] extends it to the stochastic case. Following the line of the above studies, we prove the convergence of the Eb-RHRL algorithm, i.e., $\lambda^{i}\rightarrow\lambda^{*}$ and $\tilde{u}^{i}\rightarrow\tilde{u}^{*}$ as $i\rightarrow\infty$.
@@
$\Lambda^{i+1}(\tilde{x}_{\tau})=\tilde{x}_{\tau}^{\top}Q\tilde{x}_{\tau}+\mu^{i}\left(\tilde{x}_{\tau}\right)^{\top}R\mu^{i}\left(\tilde{x}_{\tau}\right)+h({x}_{\tau})+\Lambda^{i}\left(\hat{\tilde{x}}_{\tau+1}\right),$
@@
$V^{i+1}\left(\tilde{x}_{\tau}\right)=\tilde{x}_{\tau}^{\top}Q\tilde{x}_{\tau}+\tilde{u}^{i}\left(\tilde{x}_{\tau}\right)^{\top}R\tilde{u}^{i}\left(\tilde{x}_{\tau}\right)+h({x}_{\tau})+V^{i}\left(\hat{\tilde{x}}_{\tau+1}\right),$
where $\tilde{u}^{i}\left(\tilde{x}_{\tau}\right)$ is defined by (34). We can conclude that, if $V^{0}(\cdot)=\Lambda^{0}(\cdot)=0$, then $V^{i}(\tilde{x}_{\tau})\leq\Lambda^{i}(\tilde{x}_{\tau})$ for all $i$.
@@
The control sequence $\{\tilde{u}^{i}\}$ is used to minimize $V^{i}\left(\tilde{x}_{\tau}\right)$, while all controls in $\{\mu^{i}\}$ are randoms. Thus we have $V^{i}\left(\tilde{x}_{\tau}\right)\leq\Lambda^{i}(\tilde{x}_{\tau})$. Under the condition of $V^{0}(\cdot)=\Lambda^{0}(\cdot)=0$, $V^{i}(\tilde{x}_{\tau})\leq\Lambda^{i}(\tilde{x}_{\tau})$ holds, for all $i$. ∎
Lemma 1 shows that applying $\{\tilde{u}^{i}\}$ to the robot system has a lower value function than the random control sequence. It is now ready to assist in obtaining and proving Lemma 2.
@@
Let the sequence $V^{i}\left(\tilde{x}_{\tau}\right)$ be defined as in (28). By applying an initial admissible control policy, the following conclusions hold: 1) There exists an upper bound $\bar{Z}(\tilde{x}_{\tau})$ such that $0\leq V^{i}(\tilde{x}_{\tau})\leq\bar{Z}(\tilde{x}_{\tau})$ for all $i$. 2) If the optimal control problem (27) is solvable, then there exists a upper bound $\bar{Z}(\tilde{x}_{\tau})$ such that $V(\tilde{x}_{\tau})\leq\bar{Z}(\tilde{x}_{\tau})$, where $V(\tilde{x}_{\tau})$ solves (26) and $0\leq V^{i}(\tilde{x}_{\tau})\leq V^{*}(\tilde{x}_{\tau})\leq\bar{Z}(\tilde{x}_{\tau})$ for all $i$.
@@
To begin with, $\eta(\cdot):\mathbb{R}^{n_{\mathcal{K}}}\rightarrow\mathbb{R}^{n_{u}}$ is assumed to be an admissible control policy. Let $V^{0}\left(\tilde{x}_{\tau}\right)=Z^{0}\left(\tilde{x}_{\tau}\right)=0$, where
$Z^{i+1}(\tilde{x}_{\tau})=\mathbb{E}\left[L\left(\tilde{x}_{\tau},\eta(\tilde{x}_{\tau})\right)+Z^{i}(\hat{\tilde{x}}_{\tau+1})\right]$
with $Z^{k+N}(\tilde{x}_{\tau})=\mathbb{E}\left(F(\hat{\tilde{x}}_{k+N}\right)$, for all $i\geq 1$. We define an upper bound
$\bar{Z}\left(\tilde{x}_{\tau}\right)=\mathbb{E}\left[\sum_{j=\tau}^{k+N-1}L(\tilde{x}_{j},\eta_{j}(\tilde{x}_{j}))+F(\hat{\tilde{x}}_{k+N})\right],$
with the dynamic model: $\hat{\tilde{x}}_{\tau+1}=A_{d,\tau}\tilde{x}_{\tau}+B_{d,\tau}\eta\left(\tilde{x}_{\tau}\right)$, where $L(\tilde{x}_{\tau})=\tilde{x}_{\tau}^{\top}Q\tilde{x}_{\tau}+\eta^{\top}\left(\tilde{x}_{\tau}\right)R\eta\left(\tilde{x}_{\tau}\right)+h(x_{\tau})$. With (28), the following two cases are analyzed, respectively.
@@
$Z^{i+1}(\tilde{x}_{\tau})$ $=\mathbb{E}\left[L\left(\tilde{x}_{\tau},\eta(\tilde{x}_{\tau})\right)+Z^{i}(\hat{\tilde{x}}_{\tau+1})\right]$ (39)
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+1}L\left(\tilde{x}_{j},\eta_{j}(\tilde{x}_{j})\right)+Z^{i-1}(\hat{\tilde{x}}_{\tau+2})\right]$
@@
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+i}L\left(\tilde{x}_{j},\eta_{j}(\tilde{x}_{j})\right)\right]$
Since $h(\tilde{x}_{j})\in\{0,\mu_{o}\}$, we have $L\left(\tilde{x}_{j},\eta_{j}(\tilde{x}_{j})\right)\geq 0$. For $i_{1},i_{2}\in[k,k+N-1]$, one has
$\sum_{j=\tau}^{\tau+i_{1}}L\left(x_{j},\eta_{j}(x_{j})\right)\leq\sum_{j=\tau}^{\tau+i_{2}}L\left(x_{j},\eta_{j}(x_{j})\right),i_{1}\leq i_{2}.$ (40)
@@
$Z^{i+1}(\tilde{x}_{\tau})$ (41)
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+i}L\left(\tilde{x}_{j},\eta_{j}(\tilde{x}_{j})\right)\right]$
$\leq\mathbb{E}\left[\sum_{j=\tau}^{k+N-1}L\left(\tilde{x}_{j},\eta_{j}(\tilde{x}_{j})\right)+F(\hat{\tilde{x}}_{k+N})\right]$
$=\bar{Z}(\tilde{x}_{\tau}).$
@@
$Z^{i+1}(\tilde{x}_{\tau})$ (42)
$=\mathbb{E}\left[\sum_{j=\tau}^{k+N-1}L\left(\tilde{x}_{j},\eta_{j}(\tilde{x}_{j})\right)+Z^{i-(k+N-1-\tau)}(\hat{\tilde{x}}_{k+N})\right]$
$=\mathbb{E}\left[\sum_{j=\tau}^{k+N-1}L\left(\tilde{x}_{j},\eta_{j}(\tilde{x}_{j})\right)+F(\hat{\tilde{x}}_{k+N})\right]$
$=\bar{Z}(\tilde{x}_{\tau}).$
Combining Eqs. (41) and (42), $Z^{i+1}(\tilde{x}_{\tau})\leq\bar{Z}(\tilde{x}_{\tau})$ holds.
According to Lemma 1, setting the control sequence $\{\mu^{i}\}$ as $\{\eta^{i}\}$ and the value function $\Lambda^{i}$ as $Z^{i}$, it follows that $V^{i}(\tilde{x}_{\tau})\leq Z^{i}(\tilde{x}_{\tau})\leq\bar{Z}(\tilde{x}_{\tau})$ for all $i$, which proves 1). By setting $\eta=\tilde{u}^{*}$, $V^{i}(\tilde{x}_{\tau})\leq V^{*}(\tilde{x}_{\tau})\leq\bar{Z}(\tilde{x}_{\tau})$ holds, which proves part 2). ∎
@@
According to Lemma 2, as $i\rightarrow\infty$, $V^{i}(\tilde{x}_{\tau})\rightarrow V^{*}(\tilde{x}_{\tau})$ and $\tilde{u}^{i}(\tilde{x}_{\tau})\rightarrow\tilde{u}^{*}(\tilde{x}_{\tau})$ hold. Namely, $\lambda^{i}(\tilde{x}_{\tau})$ in Eq. (35) will converge to $\lambda^{*}(\tilde{x}_{\tau})$ in Eq. (31). ∎
@@ VI Simulation and Experimental Results
In this section, we validate our proposed VF-LPC approach (i.e., Algorithm 1) on intelligent vehicles. The algorithm was deployed on a computer running Windows 11 with an Intel Core i7-11800H @2.30GHZ CPU. In the testing scenarios, we compare VF-LPC with several advanced approaches to demonstrate the superiority of VF-LPC. To further verify the effectiveness, we also tested the VF-LPC approach on an actual vehicle of Hongqi EHS-3, shown in the right part of Fig. 4. See the accompanying video for details of the simulation and experiments in this section.
@@ VI-A Planning in an Obstacle-Dense Environment
To evaluate the effectiveness of the VF-LPC approach, we compare it with advanced planning approaches in a scenario containing static obstacles; see Fig. 7. We compare VF-LPC with the commonly used planning approaches, such as Hybrid A[42], Timed Elastic Bands (TEB)[43] and OBTPAP[44] which iteratively solves the time-optimal motion planning problem. In these tests, we use the analytic kinematic vehicle model with the state variable $[X,Y,\psi]$, representing the $X$-, $Y$-coordinates and the heading, respectively. All the approaches are performed in MATLAB 2023a.
Figure 7: The planned paths generated by different comparison methods. The gray dotted line is the desired path $\mathcal{P}$ of VF-LPC.
@@ VI-A1 Evaluation metrics
To evaluate the planning methods, we employ various performance metrics. The first metric is the path length from the initial point to the destination, denoted as $\rm{Length}$; i.e., ${\rm{Length}}=\sum_{k=1}^{M-1}\lVert(X_{k+1},Y_{k+1})-(X_{k},Y_{k})\rVert$, where $M$ is the number of path points. The second metric is Aver. C.T. (average completion time of path planning). The third metric pertains to $\mathrm{safety}$, serving to quantify whether the planned paths meet the collision-free constraint.
@@ VI-A2 Analyses of the results
TABLE I: Performance Evaluations of Fig. 7
Quantitative Metrics VF-LPC TEB OBTPAP Hybrid A
$\rm{Length}$ (m) $47.54$   $47.34$ $48.44$ 46.16
Aver. C.T. (s)    $\mathbf{0.29}$      $4.80$    $5.72$   $2.29$
Safety $\checkmark$ $\checkmark$ $\checkmark$ $\times$
@@
As shown from TABLE I, the paths planned by methods VF-LPC and TEB exhibit similar lengths, albeit shorter than OBTPAP. Though the path length of hybrid A is the shortest, it is deemed unsafe due to the lack of maintaining a safe distance while avoiding the first obstacle. Note that VF-LPC is a tracking-avoidance method, while other methods could not demonstrate effectiveness in completing kinodynamic tracking control tasks. Regarding computational efficiency, VF-LPC demonstrates the lowest completion time, measured at $0.29$ s, which is significantly lower than those for hybrid A, TEB, and OBTPAP.
@@ VI-B Collision-Avoidance Simulation in High-Fidelity CarSim
Tracking a desired path while avoiding static and moving obstacles is a fundamental task. To demonstrate the superiority of VF-LPC, we compare it with MPC-CBF[45] and LMPCC[46] under different metrics in MATLAB.
@@ VI-B1 Simulation settings
$\left[\begin{array}[]{l}\dot{X}\\ \dot{Y}\\ \dot{\psi}\\ \dot{v_{x}}\\ \dot{v_{y}}\\ \dot{\omega}\\ \end{array}\right]=\left[\begin{array}[]{c}v_{x}\cos\psi-v_{y}\sin\psi\\ v_{x}\sin\psi+v_{y}\cos\psi\\ \omega\\ v_{y}\omega+a_{x}\\ 2C_{af}(\frac{\delta_{f}}{m}-\frac{v_{y}+l_{f}\omega}{mv_{x}})+2C_{ar}\frac{l_{r}\omega-v_{y}}{mv_{x}}-v_{x}\omega\\ \frac{2}{I_{z}}\left[l_{f}C_{af}(\delta_{f}-\frac{v_{y}+l_{f}\omega}{v_{x}})-l_{r}C_{ar}\frac{l_{r}\omega-v_{y}}{v_{x}}\right]\\ \end{array}\right],$ (43)
@@ VI-B2 Evaluation metrics
The desired speed $v_{d}$ of all the methods is set to $25$ km/h. To evaluate the VF-LPC approach, we compare it with other safe motion control algorithms under several metrics, which are Aver. S.T. (average solution time of each time step), $J_{\text{Lat}}$ (cost of the lateral error), $J_{\text{Heading}}$ (cost of the heading error), and $J_{\text{Con}}$ (control cost). We define a weighted average metric to evaluate the overall performance, which is
$J_{\rm{MC}}=\begin{array}[]{c}\frac{1}{M}\sum_{k=1}^{M}({Q}_{1}J_{k,\rm{Lat}}+{Q}_{2}J_{k,\rm{Heading}}+J_{k,\rm{Con}}),\end{array}$
where $J_{\rm{Lat}}=\lVert e_{y}\rVert^{2}$, $J_{\rm{Heading}}=\lVert e_{\psi}\rVert^{2}$, $J_{\rm{Con}}=\tilde{u}^{\top}{R}\tilde{u}$ is the control cost, and $M$ is the number of waypoints of the planned path. Another two quantitative metrics are the overall path length ${\rm{Length}}=\sum_{k=1}^{M-1}\lVert(X_{k+1},Y_{k+1})-(X_{k},Y_{k})\rVert$ and the completion time taken by different methods, i.e., $\rm{CT}$.

### v2

@@ V Theoretic Analysis
This section presents the theoretical analyses regarding Theorem 1. We prove that the Eb-RHRL algorithm, integrating the game-based barrier function, converges to the optimal value function and solution, thereby solving the OMP problem. Proving the convergence of the Eb-RHRL algorithm is equivalent to prove $\lambda^{i}\rightarrow\lambda^{*}$ and $\tilde{u}^{i}\rightarrow\tilde{u}^{*}$ as $i\rightarrow\infty$.
@@
$\Lambda^{i+1}(\hat{\tilde{x}}_{\tau})=\hat{\tilde{x}}_{\tau}^{\top}Q\hat{\tilde{x}}_{\tau}+\mu^{i}(\hat{\tilde{x}}_{\tau})^{\top}R\mu^{i}(\hat{\tilde{x}}_{\tau})+h(\hat{x}_{\tau})+\Lambda^{i}(\hat{\tilde{x}}_{\tau+1}),$
@@
$V^{i+1}(\hat{\tilde{x}}_{\tau})=\hat{\tilde{x}}_{\tau}^{\top}Q\hat{\tilde{x}}_{\tau}+\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})^{\top}R\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})+h(\hat{x}_{\tau})+V^{i}(\hat{\tilde{x}}_{\tau+1}),$
where $\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})$ is defined by (34). We can conclude that, if $V^{0}(\cdot)=\Lambda^{0}(\cdot)=0$, then $V^{i}(\hat{\tilde{x}}_{\tau})\leq\Lambda^{i}(\hat{\tilde{x}}_{\tau})$ for all $i$.
@@
[…]
Lemma 1 shows that applying $\{\tilde{u}^{i}\}$ to the robot system results in a lower value function than applying a random solution sequence. It is now used to obtain and prove Lemma 2.
@@
Let the sequence $V^{i}(\hat{\tilde{x}}_{\tau})$ be defined as in (28). By applying an initial admissible control policy, the following conclusions hold: 1) There exists an upper bound $\bar{Z}(\hat{\tilde{x}}_{\tau})$ such that $0\leq V^{i}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}(\hat{\tilde{x}}_{\tau})$ for all $i$. 2) If (27) is solvable, then there exists a upper bound $\bar{Z}(\hat{\tilde{x}}_{\tau})$ such that $V(\hat{\tilde{x}}_{\tau})\leq\bar{Z}(\hat{\tilde{x}}_{\tau})$, where $V(\hat{\tilde{x}}_{\tau})$ solves (26) and $0\leq V^{i}(\hat{\tilde{x}}_{\tau})\leq V^{*}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}(\hat{\tilde{x}}_{\tau})$ for all $i$.
@@
To begin with, $\eta(\cdot):\mathbb{R}^{n_{\mathcal{K}}}\rightarrow\mathbb{R}^{n_{u}}$ is assumed to be an admissible policy. Let $V^{0}(\hat{\tilde{x}}_{\tau})=Z^{0}(\hat{\tilde{x}}_{\tau})=0$, where
$Z^{i+1}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\eta(\hat{\tilde{x}}_{\tau}))+Z^{i}(\hat{\tilde{x}}_{\tau+1})\right]$
with $Z^{k+N}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left(F(\hat{\tilde{x}}_{k+N}\right)$, for all $i\geq 1$. We define an upper bound
$\bar{Z}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[\sum_{j=\tau}^{k+N-1}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))+F(\hat{\tilde{x}}_{k+N})\right],$
with the dynamic model: $\hat{\tilde{x}}_{\tau+1}=A_{d,\tau}\hat{\tilde{x}}_{\tau}+B_{d,\tau}\eta(\hat{\tilde{x}}_{\tau})$, where $L(\hat{\tilde{x}}_{\tau})=\hat{\tilde{x}}_{\tau}^{\top}Q\hat{\tilde{x}}_{\tau}+\eta^{\top}(\hat{\tilde{x}}_{\tau})R\eta(\hat{\tilde{x}}_{\tau})+h(\hat{x}_{\tau})$. With (28), the following two cases are analyzed, respectively.
@@
$Z^{i+1}(\hat{\tilde{x}}_{\tau})$ $=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\eta(\hat{\tilde{x}}_{\tau}))+Z^{i}(\hat{\tilde{x}}_{\tau+1})\right]$ (39)
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+1}L\left(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j})\right)+Z^{i-1}(\hat{\tilde{x}}_{\tau+2})\right]$
@@
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+i}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))\right]$
Since $h(\hat{x}_{j})\in\{0,\mu_{o}\}$, we have $L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))\geq 0$. For $i_{1},i_{2}\in[k,k+N-1]$, one has
$\sum_{j=\tau}^{\tau+i_{1}}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))\leq\sum_{j=\tau}^{\tau+i_{2}}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j})),i_{1}\leq i_{2}.$ (40)
@@
$Z^{i+1}(\hat{\tilde{x}}_{\tau})$ (41)
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+i}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))\right]$
$\leq\mathbb{E}\left[\sum_{j=\tau}^{k+N-1}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))+F(\hat{\tilde{x}}_{k+N})\right]=\bar{Z}(\hat{\tilde{x}}_{\tau}).$
@@
[…]
Combining Eqs. (41) and (42), $Z^{i+1}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}(\hat{\tilde{x}}_{\tau})$ holds.
[…]
@@
[…]
@@ VI Simulation and Experimental Results
In this section, we present the simulation and experimental results to validate our proposed VF-LPC approach. The algorithm was deployed on a computer running Windows 11 with an Intel Core i7-11800H @2.30GHZ CPU. We first compared VF-LPC with advanced IMPC methods in CarSim software. Then the effectiveness of the online model compensating strategy is verified through a tracking control task. The simulation results on a quadrotor UAV were performed. We also tested the VF-LPC approach on an actual ground vehicle of Hongqi EHS-3. The platform is shown in the right part of Fig. 4.
@@ VI-A Planning in an Obstacle-Dense Environment
To evaluate the effectiveness of the VF-LPC approach, we compare it with advanced planning approaches in a scenario containing static obstacles; see Fig. 7. We compare VF-LPC with the commonly used planning approaches, such as Hybrid A*[38], RRT*-CFS[39], Timed Elastic Bands (TEB)[40] and OBTPAP[41] which iteratively solves the time-optimal motion planning problem. In these tests, we use the analytic kinematic vehicle model with the state variable $[X,Y,\psi]$, representing the $X$-, $Y$-coordinates and the heading, respectively. All the approaches are performed in MATLAB 2023a.
@@ VI-A1 Evaluation metrics
To evaluate the planning methods, we employ various performance metrics. The first metric is the trajectory length from the initial point to the destination, denoted as $L_{p}$; i.e., ${L}_{p}=\sum_{k=1}^{M-1}\lVert(X_{k+1},Y_{k+1})-(X_{k},Y_{k})\rVert$, where $M$ is the number of trajectory points. The second metric $t_{\mathrm{cal}}$ is the $\mathrm{calculation\ time}$ of trajectory planning. The third metric pertains to $\mathrm{SD}$, serving to quantify whether the planned trajectories have a safe distance from obstacles.
@@ VI-A2 Analyses of the results
As shown from TABLE I, the paths planned by methods VF-LPC and TEB exhibit similar lengths, albeit shorter than OBTPAP. Though the trajectory length of Hybrid A* is the shortest, it is deemed unsafe due to the lack of maintaining a safe distance while avoiding the first obstacle. Regarding computational efficiency, VF-LPC demonstrates the best performance among the methods.
TABLE I: The Comparative Analysis of Planning Results using Different Methods in the Obstacle-dense Environment.
Metrics VF-LPC Hybrid A*[38] RRT*-CFS[39] TEB[40] OBTPAP[41]
${L}_{p}$ (m)   $47.44$   46.16 $54.21$ $48.34$ $47.54$
$t_{\mathrm{cal}}$ (s)     $\mathbf{0.29}$    $2.29$ $50.28$    $4.80$    $5.72$
$\mathrm{SD}$? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Figure 7: The planned results using different comparison methods.
@@ VI-B Collision-Avoidance Simulation in High-Fidelity CarSim
Tracking a desired path while avoiding static and moving obstacles is a fundamental task. To demonstrate the superiority of VF-LPC, we compare it with MPC-CBF[42], LMPCC[43], RHRL-KDP[10], and CFS[44] under different metrics.
@@ VI-B1 Simulation settings
$\left[\begin{array}[]{l}\dot{X}\\ \dot{Y}\\ \dot{\psi}\\ \dot{v}_{x}\\ \dot{v}_{y}\\ \dot{\omega}\\ \end{array}\right]=\left[\begin{array}[]{c}v_{x}\cos\psi-v_{y}\sin\psi\\ v_{x}\sin\psi+v_{y}\cos\psi\\ \omega\\ v_{y}\omega+a_{x}\\ 2C_{af}(\frac{\delta_{f}}{m}-\frac{v_{y}+l_{f}\omega}{mv_{x}})+2C_{ar}\frac{l_{r}\omega-v_{y}}{mv_{x}}-v_{x}\omega\\ \frac{2}{I_{z}}\left[l_{f}C_{af}(\delta_{f}-\frac{v_{y}+l_{f}\omega}{v_{x}})-l_{r}C_{ar}\frac{l_{r}\omega-v_{y}}{v_{x}}\right]\\ \end{array}\right],$
@@ VI-B2 Evaluation metrics
The desired speed $v_{d}$ of all the methods is set to $25$ km/h. To evaluate the VF-LPC approach, we compare it with other IMPC algorithms under several metrics, which are Aver. S.T. (average solution time of each time step), $J_{\rm{Lat}}$ (cost of the lateral error), $J_{\rm{Heading}}$ (cost of the heading error), and $J_{\rm{Con}}$ (control cost). Specifically, they are defined by $J_{\rm{Lat}}=\lVert(X-X_{r})\sin\psi_{r}-(Y-Y_{r})\cos\psi_{r}\rVert_{q_{1}}^{2}$, $J_{\rm{Heading}}=\lVert\psi-\psi_{r}\rVert_{q_{2}}^{2}$, and $J_{\rm{Con}}=\lVert\tilde{u}\rVert_{R}^{2}$, where $q_{1},q_{2}>0$ and $R\in\mathbb{R}^{n_{u}\times n_{u}}$ is definite. We define a weighted average cost to evaluate the overall performance, which is
$J_{\rm{MC}}=\begin{array}[]{c}\frac{1}{M}\sum_{k=1}^{M}(J_{k,\rm{Lat}}+J_{k,\rm{Heading}}+J_{k,\rm{Con}}),\end{array}$
where $M$ is the number of waypoints of the driving trajectory. The other quantitative metrics are route length $L_{p}$, completion time $t_{\mathrm{com}}$, and calculation time $t_{\rm{cal}}$.